OpenStudy (anonymous):

The width of a rectangular poster is half its length. The area of the poster is 128 square feet. What is the width of the poster?

OpenStudy (anonymous):

area of rectangle = L*W =128 L* (0.5)L=128 L*L=256 root 256=16 L=16 W=0.5L=8

OpenStudy (cwrw238):